OpenStudy (anonymous):

What are the factors of –56 that would be relevant in the correct factorization of x^2 - x - 56? A. (–7, –8) B. (–7, 8) C. (7, –8) D. (7, 8)

OpenStudy (anonymous):

@hockeychick23 @samjordon

OpenStudy (anonymous):

so find 2 numbers that when you add them equal -1 and when you multiply them equal -56 So lets list the factors of 56 1*56=56 2*28=56 4*18=56 7*8=56

OpenStudy (anonymous):

So which 2 factors when you multiply them equals -56 and when you add them equals -1 Please keep in mind that the factors can be negative or positive So for instance if you choose the factors 7 and 8 obviously either 7 or 8 will be negative

OpenStudy (campbell_st):

well you look at a question of this type... the sign of the middle term is critical in you question the factors of -56 add to -1 this means that the larger factor is negative... a negative constant and negative middle -x indicate that there is 1 positive and 1 negative factor... if you have \[x^2 + x - 56\] the factors of -56 that add to 1 the larger factor is positive in this case... hope it helps for your future work...